# A new intersection condition in extremal set theory

### v1

@@ 4.1. General observations in the non-uniform case
Let $\mathcal{F}\subset 2^{n}$, $F\subset G$, $F\in\mathcal{F}$ and $G\notin\mathcal{F}$. We call $\mathcal{F}-\{F\}\cup\{G\}$ as an upward-shifting of $\mathcal{F}$. A family $\mathcal{F^{\prime}}$ is called an up-shifted version of $\mathcal{F}$ if $\mathcal{F^{\prime}}$ can be obtained from $\mathcal{F}$ by some upward-shifting.
@@ 4.2. Sets with large intersection
Let $\overline{d(\mathcal{A})}=2f(\mathcal{A},1)+3f(\mathcal{A},2)+3f(\mathcal{A},3)$ and $\overline{d(A,B)}=2f(\{A,B\},1)+3f(\{A,B\},2)$. Denote by $\overline{\mathbf{H}_{\ell}}$ the class of families such that $\mathcal{H}\in\overline{\mathbf{H}_{\ell}}$ and $\mathcal{A}\subset\mathcal{H}$ imply $\overline{d(\mathcal{A})}\leq\ell$. Let =$\mathcal{F}^{-}=\{\overline{F}:F\in\mathcal{F}\}$ and let $\overline{h}(n,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset 2^{[n]},\mathcal{F}\in\overline{\mathbf{H}_{\ell}}\}$.
@@
The following proof is not working when $p=0$. It is easy to see that in this case $h(n,3n-x)=2$ if $0\leq x\leq 3$ and $h(n,3n-x)=3$ if $4\leq x\leq 5$.
@@
Let =$\mathcal{F}=\{F:|F|\leq p\}$ be a family. By Claim 4.7, for all $\mathcal{A}\subset\mathcal{F}$, $\overline{d(\mathcal{A})}\leq 2(|A|+|B|+|C|)=6p$, which means $\mathcal{F}\in\overline{\mathbf{H}_{6p}}\subset\overline{\mathbf{H}_{x}}$. ∎
@@
Denote the family ($\mathcal{F}\cap{[n]\choose i}$ by $\mathcal{F}_{i}$.
@@
Case 1: $q=0$ or $q=1$. By Claim 4.11, if $|\mathcal{F}|=\overline{h}(n,x)$, then all sets in $\mathcal{F}$ have at most $p$ elements. From here it is easy to see that $|\mathcal{F}|\leq\sum_{i=0}^{p}{n\choose p}$ and =$\mathcal{F}=\{F:|F|\leq p\}$ is a maximal family.
@@
If there exists $A\in\mathcal{F}$ such that $|A|=p+2$, then $|A\cap B|\geq 4$ for all $B\in\mathcal{F}_{p+2}$ and $|A\cap C|\geq 2$ for all $C\in\mathcal{F}_{p+1}$ by Claim 4.10. So in this case $|\mathcal{F}_{p+1}|\leq{p+2\choose 2}{n-2\choose p-1}$ and $|\mathcal{F}_{p+2}|\leq{n-4\choose p-2}$ by Theorem 1.6. Hence we have ||($|\mathcal{F}|\leq{n-4\choose p-2}+{p+2\choose 2}{n-2\choose p-1}+\sum_{i=0}^{p}{n\choose p}$.
If there are no sets with $p+2$ elements, then all sets have at most $p+1$ elements. If there exist sets $A$, $B$, $C\in\mathcal{F}_{p+1}$ satisfying $d(\mathcal{A})\leq 1$, then $\overline{d(\mathcal{A})}\geq 6p+5$ by Claim 4.7. So $\mathcal{F}_{p+1}\in\mathbf{H}_{2}$. Since $n\geq 2^{3p+2}p^{2}+p+1\geq 4(p+1)^{3}$, we can use Theorem 2.5, so $\mathcal{F}$ contains at most (${n-1\choose p}$ sets with $p+1$ elements so ||($|\mathcal{F}|\leq{n-1\choose p}+\sum_{i=0}^{p}{n\choose p}$.
[…]
Case 4: $q=3$. Since =$\mathcal{F}=\{F:|F|\leq p$ or ($|F|=p+1$ and $1\in F)\}\in\overline{\mathbf{H}_{6p+3}}$, then (${n-1\choose p}+\sum_{i=0}^{p}{n\choose p}\leq\overline{h}(n,6p+3)$.
@@
[…]

### v2

@@ 4.1. General observations in the non-uniform case
Let $\mathcal{F}\subset 2^{n}$, $\subset G, FF\in\mathcal{F}$ and $G\notin\mathcal{F}$. We call $\mathcal{F}-\{F\}\cup\{G\}$ as an upward-shifting of $\mathcal{F}$. A family $\mathcal{F^{\prime}}$ is called an up-shifted version of $\mathcal{F}$ if $\mathcal{F^{\prime}}$ can be obtained from $\mathcal{F}$ by some upward-shifting.
@@ 4.2. Sets with large intersection
Let $\overline{d(\mathcal{A})}=2f(\mathcal{A},1)+3f(\mathcal{A},2)+3f(\mathcal{A},3)$ and $\overline{d(A,B)}=2f(\{A,B\},1)+3f(\{A,B\},2)$. Denote by $\overline{\mathbf{H}_{\ell}}$ the class of families such that $\mathcal{H}\in\overline{\mathbf{H}_{\ell}}$ and $\mathcal{A}\subset\mathcal{H}$ imply $\overline{d(\mathcal{A})}\leq\ell$. Let $\mathcal{F}^{-}=\{\overline{F}:F\in\mathcal{F}\}$ and let $\overline{h}(n,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset 2^{[n]},\mathcal{F}\in\overline{\mathbf{H}_{\ell}}\}$.
@@
The following proof is not working when $p=0$. It is easy to see that in this case h(n,3nx)=2h(n,3n-x)=2 if $0\leq x\leq 3$ and h(n,3nx)=3h(n,3n-x)=3 if $4\leq x\leq 5$.
@@
Let $\mathcal{F}=\{F:|F|\leq p\}$ be a family. By Claim 4.7, for all $\mathcal{A}\subset\mathcal{F}$, $\overline{d(\mathcal{A})}\leq 2(|A|+|B|+|C|)=6p$, which means $\mathcal{F}\in\overline{\mathbf{H}_{6p}}\subset\overline{\mathbf{H}_{x}}$. ∎
@@
Denote the family $\mathcal{F}\cap{[n]\choose i}$ by $\mathcal{F}_{i}$.
@@
Case 1: $q=0$ or $q=1$. By Claim 4.11, if $|\mathcal{F}|=\overline{h}(n,x)$, then all sets in $\mathcal{F}$ have at most $p$ elements. From here it is easy to see that $|\mathcal{F}|\leq\sum_{i=0}^{p}{n\choose p}$ and $\mathcal{F}=\{F:|F|\leq p\}$ is a maximal family.
@@
If there exists $A\in\mathcal{F}$ such that $|A|=p+2$, then $|A\cap B|\geq 4$ for all $B\in\mathcal{F}_{p+2}$ and $|A\cap C|\geq 2$ for all $C\in\mathcal{F}_{p+1}$ by Claim 4.10. So in this case $|\mathcal{F}_{p+1}|\leq{p+2\choose 2}{n-2\choose p-1}$ and $|\mathcal{F}_{p+2}|\leq{n-4\choose p-2}$ by Theorem 1.6. Hence we have $|\mathcal{F}|\leq{n-4\choose p-2}+{p+2\choose 2}{n-2\choose p-1}+\sum_{i=0}^{p}{n\choose p}$.
If there are no sets with $p+2$ elements, then all sets have at most $p+1$ elements. If there exist sets $A$, $B$, $C\in\mathcal{F}_{p+1}$ satisfying $d(\mathcal{A})\leq 1$, then $\overline{d(\mathcal{A})}\geq 6p+5$ by Claim 4.7. So $\mathcal{F}_{p+1}\in\mathbf{H}_{2}$. Since $n\geq 2^{3p+2}p^{2}+p+1\geq 4(p+1)^{3}$, we can use Theorem 2.5, so $\mathcal{F}$ contains at most $n-1\choose p$} sets with $p+1$ elements so $|\mathcal{F}|\leq{n-1\choose p}+\sum_{i=0}^{p}{n\choose p}$.
Comparing the two cases, we can see that the upper bound obtained in the latter case is larger, so $|\mathcal{F}|\leq{n-1\choose p}+\sum_{i=0}^{p}{n\choose p}$ and equality holds if $\mathcal{F}=\{F:|F|\leq p$ or $(|F|=p+1$ and $1\in F)\}$.
Case 4: $q=3$. Since $\mathcal{F}=\{F:|F|\leq p$ or ($|F|=p+1$ and $1\in F)\}\in\overline{\mathbf{H}_{6p+3}}$, then ${n-1\choose p}+\sum_{i=0}^{p}{n\choose p}\leq\overline{h}(n,6p+3)$.
@@
Summarizing (1), (2), (3): $|\mathcal{F}|\leq\max\{{n-3\choose p-1}+{2p+3\choose 2}{n\choose p-1},{n-3\choose p-1}+{p+2\choose 2}{n\choose p-1},{n-1\choose p}+{n-2\choose p}\}+\sum_{i=0}^{p}{n\choose p}$. It can be proven by calculation that ${n-3\choose p-1}\leq{n-2\choose p}$ and ${p+2\choose 2}{n\choose p-1}\leq{2p+3\choose 2}{n\choose p-1}\leq{n-1\choose p}$ if $n\geq 8p^{3}$, so $|\mathcal{F}|\leq{n-1\choose p}+{n-2\choose p}+\sum_{i=0}^{p}{n\choose p}$. Equality holds if $\mathcal{F}=\{F:|F|\leq p$ or $(|F|=p+1$ and $|\{1,2\}\cap F|\geq 1)\}$. $\square$